# Re-anchoring Quantum Monte Carlo with Tensor-Train Sketching

### v4

@@ Appendix B Details on Convergence Analysis of cp-AFQMC
Hamiltonians $H_{1}$ and $H_{2}$; initial walkers $\widehat{\Psi}^{(0)}=\sum_{k}w_{k}^{(0)}\frac{\phi_{k}^{(0)}}{O_{\mathrm{tr}}(\phi_{k}^{(0)})}$; trial wavefunction $\Psi_{\mathrm{tr}}$, step size $\Delta\tau$, number of iteration steps $M$.
Walkers $\widehat{\Psi}^{(M)}=\sum_{k}w_{k}^{(M)}\frac{\phi_{k}^{(M)}}{O_{\mathrm{tr}}(\phi_{k}^{(M)})}$
@@
      Let $\phi^{(n)\prime}_{k}\leftarrow B_{H_{1}/2}\phi^{(n)}_{k}$, $w_{k}^{(n)}\leftarrow w_{k}^{(n)}\frac{O_{\mathrm{tr}}(\phi^{(n)\prime}_{k})}{O_{\mathrm{tr}}(\phi^{(n)}_{k})}$.
@@
      Update $\phi_{k}^{(n)\prime}$ to $\phi_{k}^{(n)\prime\prime}$ by applying $B_{H_{2}}$ to $\phi_{k}^{(n)\prime}$ stochastically using the decomposition in (49). Update weights $w_{k}^{(n)}$ to $w_{k}^{(n+1)}$. The details are provided in Section C.1.
@@
      Let $\phi^{(n+1)}_{k}\leftarrow B_{H_{1}/2}\phi^{(n)\prime\prime}_{k}$, $w_{k}^{(n+1)}\leftarrow w_{k}^{(n+1)}\frac{O_{\mathrm{tr}}(\phi^{(n+1)}_{k})}{O_{\mathrm{tr}}(\phi^{(n)\prime\prime}_{k})}$.
  • Periodically estimate the ground-state energy.
  • Periodically do population control.
@@
C.1. Details for applying $B_{H_{2}}$
In this section, we detail how to apply $B_{H_{2}}$ to a walker $\phi_{k}^{(n)\prime}$ stochastically. To summarize, we sample $x_{ij}$ for each $\langle i,j\rangle$ one at a time, and apply a single $b(x_{ij})$ to $\phi^{(n)\prime}_{k}$. For example, for a single pair $\langle i,j\rangle$, we let
$\sum_{x_{ij}}p(x_{ij})b(x_{ij})\frac{w_{k}^{(n)}\phi_{k}^{(n)\prime}}{O_{\mathrm{tr}}(\phi_{k}^{(n)\prime})}=\sum_{x_{ij}}\widetilde{p}_{ij,k}^{(n)}(x_{ij})w_{k}^{(n)}\frac{p(x_{ij})}{\widetilde{p}_{ij,k}^{(n)}(x_{ij})}\frac{b(x_{ij})\phi_{k}^{(n)\prime}}{O_{\mathrm{tr}}(\phi_{k}^{(n)\prime})}\\ \approx w_{k}^{(n)}\frac{p(x^{(n)}_{ij,k})}{\widetilde{p}_{ij,k}^{(n)}(x^{(n)}_{ij,k})}\frac{b(x^{(n)}_{ij,k})\phi_{k}^{(n)\prime}}{O_{\mathrm{tr}}(\phi_{k}^{(n)\prime})},$
where we obtain $x^{(n)}_{ij,k}$ by importance sampling according to a new distribution $\widetilde{p}_{ij,k}^{(n)}$. Similar to Section 2.1, we let
$\widetilde{p}_{ij,k}^{(n)}(x_{ij})=\frac{1}{\mathcal{N}^{(n)}_{ij,k}}\frac{O_{\mathrm{tr}}(b(x_{ij})\phi_{k}^{(n)\prime})}{O_{\mathrm{tr}}(\phi_{k}^{(n)\prime})}p(x_{ij}),$
where $p(x_{ij})=1/2$ for $x_{ij}=\pm 1$. In this case, the update rule for the weight and the walker becomes
$w^{(n)}_{k}\leftarrow\begin{cases}w_{k}^{(n)}\mathcal{N}^{(n)}_{ij,k}\quad&\text{if}\ \langle\Psi_{\mathrm{tr}},b(x^{(n)}_{ij,k})\phi_{k}^{(n)\prime}\rangle>0,\\ 0&\text{otherwise},\end{cases}$
@@ C.1. Details for applying $B_{H_{2}}$
$\phi^{(n)\prime\prime}_{k}=b(x^{(n)}_{ij,k})\phi_{k}^{(n)\prime}.$
This procedure is then repeated for all pairs $\langle i,j\rangle$.
Appendix D Systematic Error of Vanilla cp-AFQMC on TFI System
In this section, we study the origin of the bias in cp-AFQMC, using the TFI system with $16$ spins in 1D as an example. As mentioned earlier in Section 5, the TFI system is free of sign problem with a properly chosen trial wavefunction, but there is still a systematic error caused by the fixed imaginary-time step $\Delta\tau$ and the finite population size $N$.
At each iteration step, a Trotter error $O(\Delta\tau^{3})$ is introduced, as shown in (47). For a fixed imaginary-time, one needs to run $O(1/\Delta\tau)$ steps of random walk. Thus, the total error in terms of the energy estimation accumulates to $O(\Delta\tau^{2})$. Ideally, the Trotter error will vanish as the imaginary-time step $\Delta\tau$ goes to 00. However, this is not always the case as shown in Figure 14(a). With $2000$ walkers, the estimated energy indeed converges to the ground-state energy as $\Delta\tau$ goes to 00. If we use $200$ walkers, however, there is still a quite obvious bias even when $\Delta\tau$ becomes 00. Moreover, with fewer walkers, the bias is consistently larger with any imaginary-time step $\Delta\tau$.
@@ Appendix D Systematic Error of Vanilla cp-AFQMC on TFI System
To further study how the error depends on the number of walkers, we fix the imaginary-time step to $\Delta\tau=0.01$, and run cp-AFQMC with different population size. As shown in Figure 14(b), the error is roughly inversely proportional to the population size. The bias is introduced by the population control procedure summarized in Algorithm 5. To be more specific, if we look at each single walker, it breaks into a random number of copies, with the average number of copies matching its weight. However, to maintain a constant population size, the number of copies must be adjusted based on the total, which introduces dependency among walkers. This dependency leads to a bias in the energy estimation.
Algorithm 5 Population control
walkers $\phi_{k}$, weights $w_{k}$, and $O_{\mathrm{tr}}(\phi_{k})$.
adjusted walkers $\phi_{k}$, weights $w_{k}$, and $O_{\mathrm{tr}}(\phi_{k})$.
normalize the weights so that the average value of the weights is $1$.
$s\leftarrow\text{a uniformly random number in }(0,1)$,
$n_{1}\leftarrow 1$,
for $k=1,2,\cdots,N$ do
     $s\leftarrow s+w_{k}$;
     $n_{2}\leftarrow\text{the smallest integer that is greater than or equal to }s$;
     set the walkers $\phi_{n_{1}},\cdots,\phi_{n_{2}-1}$ to be $\phi_{k}$, and corresponding overlaps to be $O_{\mathrm{tr}}(\phi_{k})$;
     $n_{1}\leftarrow n_{2}$
end for
reset all the weights to be $w_{k}=1$.

### v5

@@ C.1. Implementation details of cp-AFQMC
Hamiltonians $H_{1}$ and $H_{2}$; initial walkers $\widehat{\Psi}^{(0)}=\sum_{k}w_{k}^{(0)}\frac{\phi_{k}^{(0)}}{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(0)}}}$; trial wavefunction $\Psi_{\mathrm{tr}}$, step size $\Delta\tau$, number of iteration steps $M$.
Walkers $\widehat{\Psi}^{(M)}=\sum_{k}w_{k}^{(M)}\frac{\phi_{k}^{(M)}}{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(M)}}}$
@@
      Let $\phi^{(n)\prime}_{k}\leftarrow B_{H_{1}/2}\phi^{(n)}_{k}$, $w_{k}^{(n)}\leftarrow w_{k}^{(n)}\frac{\max\{\braket{\Psi_{\mathrm{tr}},\phi^{(n)\prime}_{k}},0\}}{\braket{\Psi_{\mathrm{tr}},\phi^{(n)}_{k}}}$.
@@
      Update $\phi_{k}^{(n)\prime}$ to $\phi_{k}^{(n)\prime\prime}$ by applying $B_{H_{2}}$ to $\phi_{k}^{(n)\prime}$ stochastically using the decomposition in (36). Update weights $w_{k}^{(n)}$ to $w_{k}^{(n+1)}$ as in (37).
@@
      Let $\phi^{(n+1)}_{k}\leftarrow B_{H_{1}/2}\phi^{(n)\prime\prime}_{k}$, $w_{k}^{(n+1)}\leftarrow w_{k}^{(n+1)}\frac{\max\{\braket{\Psi_{\mathrm{tr}},\phi^{(n+1)}_{k}},0\}}{\braket{\Psi_{\mathrm{tr}},\phi^{(n)\prime\prime}_{k}}}$.
  • Periodically estimate the ground-state energy as in (33).
  • Periodically apply population control as in Algorithm 3.
@@
We now detail how to apply $B_{H_{2}}$ to a walker $\phi_{k}^{(n)\prime}$ stochastically. We sample $x_{ij}$ for each $\langle i,j\rangle$ one at a time, and apply a single $b(x_{ij})$ to $\phi^{(n)\prime}_{k}$. For example, for a single pair $\langle i,j\rangle$, if at least one $x_{ij}$ satisfies $\langle\Psi_{\mathrm{tr}},b(x_{ij})\phi_{k}^{(n)\prime}\rangle>0$, then we define the probability distribution function
$q_{ij,k}^{(n)}(x_{ij})=\frac{1}{\mathcal{N}^{(n)}_{ij,k}}\frac{\max\{\langle\Psi_{\mathrm{tr}},b(x_{ij})\phi_{k}^{(n)\prime}\rangle,0\}}{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)\prime}}}p(x_{ij}).$
The update rule for the weight and the walker becomes
(37) $w^{(n+1)}_{k}=w_{k}^{(n)}\mathcal{N}^{(n)}_{ij,k}$.
@@
$\phi^{(n)\prime\prime}_{k}=b(x^{(n)}_{ij,k})\phi_{k}^{(n)\prime},x_{ij,k}^{(n)}\sim q_{ij,k}^{(n)}.$
If $\langle\Psi_{\mathrm{tr}},b(x_{ij})\phi_{k}^{(n)\prime}\rangle\leq 0$ for both $x_{ij}=\pm 1$, then we remove this walker by setting $w_{k}^{(n+1)}=0$. This procedure is then repeated for all pairs $\langle i,j\rangle$.
Appendix D Systematic Error of cp-AFQMC on TFI Systems
In this section, we study the origin of the bias in cp-AFQMC, using the TFI system with $16$ spins in 1D as an example. As mentioned in Section 5, the TFI system is free of sign problem with a properly chosen trial wavefunction, but there is still a systematic error caused by the fixed imaginary-time step $\Delta\tau$ and the finite population size $N$.
At each iteration step, a Trotter error $O(\Delta\tau^{3})$ is introduced, as shown in (34). For a fixed imaginary-time, one needs to run $O(1/\Delta\tau)$ steps of random walk. Thus, the total error in terms of the energy estimation accumulates to $O(\Delta\tau^{2})$. Ideally, the Trotter error will vanish as the imaginary-time step $\Delta\tau$ goes to 0. However, this is not always the case as shown in Figure 15(a). With $2000$ walkers, the estimated energy indeed converges to the ground-state energy as $\Delta\tau$ goes to 0. If we use $200$ walkers, however, there is still a quite obvious bias even when $\Delta\tau$ becomes 0. Moreover, with fewer walkers, the bias is consistently larger with any imaginary-time step $\Delta\tau$.
@@ Appendix D Systematic Error of cp-AFQMC on TFI Systems
To further study how the error depends on the number of walkers, we fix the imaginary-time step to $\Delta\tau=0.01$, and run cp-AFQMC with different population size. As shown in Figure 15(b), the error is roughly inversely proportional to the population size. The bias is introduced by the population control procedure as in Algorithm 3. To be more specific, if we look at each single walker, it breaks into a random number of copies, with the average number of copies matching its weight. However, to maintain a constant population size, the number of copies must be adjusted based on the total, which introduces dependency among walkers. This dependency leads to a bias in the energy estimation.